OpenStudy (anonymous):

WILL FAN AND MEDAL!!!! Solve the inequality: 4x – 6 > 6x – 20

OpenStudy (anonymous):

best explanation gets a fan and medal

pooja195 (pooja195):

pretend this > is an equal sign 4x – 6 = 6x – 20 solve for x

OpenStudy (anonymous):

ik tht part but idk the rest how do u solve it

pooja195 (pooja195):

We want to isolate x right? start by adding 6 to both sides -20+6=?

OpenStudy (anonymous):

-14

pooja195 (pooja195):

good now we have 4x = 6x – 14 subtract 6x from both sides whats 4-6=?

OpenStudy (anonymous):

-2

pooja195 (pooja195):

ok -2x> -14 divide BOTH sides by -2 -2/-2=? -14/-2=?

OpenStudy (anonymous):

7

pooja195 (pooja195):

good since we divided by a NEGATIVE we need to FLIP the sign so it would be x<7

OpenStudy (anonymous):

thank you sooo much!!!!!

pooja195 (pooja195):

:)